# ”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder

### v1

@@ 4.1. Method
Table 2. Demographic information and mental health experience of the 7 OCD professionals in Study 2.
4.1.2. Procedure.
All interviews were conducted via Zoom. Each interview consists of three phases that overall lasted about 90 minutes. The first phase focused on participants’ demographic information, certificate, and experience as a mental health care provider. In the second phase, we asked about participants’ strategies and technologies used (if applicable) in different treatment stages. If a participant conducted ERP, we further asked for details about exposure creation and progress monitoring during exposure practices. We then asked about out-of-session homework and external support clients might need. In the last phase, we prompted for the challenges of current therapy practices as well as their opinions on current technological support for OCD.
4.1.3. Data Analysis.
Similar to Study 1, we conducted thematic analysis on interview transcripts. Two researchers coded two representative transcripts (samples) independently and generated a initial codebook. One researcher coded the remaining five transcripts using the codebook, updating it with new codes after discussion and agreement with the other researcher. Following a hybrid approach (Swain, 2018; Braun and Clarke, 2006), we first generated themes deductively based on therapy treatment process and challenges identified in prior work (e.g., psychoeducation) (Blakey et al., 2017; Wu and Lewin, 2017; Yule and Whittal, 2017). We then generate sub-themes and themes (e.g. using relaxation techniques suitably) inductively (Boyatzis, 1998). The final codebook contained over 300 codes, which we grouped into 16 themes and sub-themes (Table A.3).
4.2. Findings
4.2.1. OCD Diagnosis and Assessment
In the initial assessment of a client’s OCD symptoms, all participants used standardized self symptom rating tools, such as Y-BOCS (Goodman et al., 1989), to gain a basic understanding of their obsessions and compulsions. Since compulsions can often manifest in ways clients might not even realize, such as trigger avoidance and mental compulsion (e.g., rumination), our therapists usually conducted interviews to identify the presence of such mental compulsions (T1-T3, T5, T7). Due to clients’ lack of awareness of avoidance or mental compulsions, identifying them in early stage is difficult but important for treatment.
Participants developed criteria to identify compulsive behaviors and assess their impact on daily life:
@@ 4.2.1. OCD Diagnosis and Assessment
In addition to in-session assessment, homework was assigned for clients to self-monitor their OCD symptoms now that they have better awareness of their OCD (T1-T3, T5-T7). Overall, these strategies could potentially be integrated into self-management technologies to help improve OCD people’s symptoms self-awareness.
4.2.2. Psychoeducation to Gain Insights
A person’s insight in OCD refers to the degree to which one recognizes the irrational and excessive nature of their OCD symptoms (Wu and Lewin, 2017), which is considered to be a key factors for successful treatment (T1, T4, T6, T7). Clients usually gain insights through psychoeducation along with the initial assessment, where they were taught about the definition of OCD, how OCD therapy works, the expected outcome of the therapy (T1-T7). Psychoeducation happens not only didactically but also throughout the treatment process (T1-T4, T6, T7). Clients gradually gain insights by discussing their symptoms during therapy sessions and symptom self-tracking between sessions.
Moreover, to assist with in-session treatment, most participants believed that psychoeducation for clients’ family is equally necessary (T1-5) because clients’ family need to learn how to respond to their loved ones’ behaviors in a therapeutic and helpful way instead of accommodating their OCD (e.g., providing reassurance) (T2-4).
As such, psychoeducation should also be considered to incorporate into self-management technology for both OCD clients and their families to enhance insights and awareness.
4.2.3. Active Treatment via ERP
All participants conducted ERP for their clients. In this section, we reported participants’ strategies on identifying trigger intensity, creating exposures, and guiding clients through exposure practices while monitoring their symptoms and progress.
Building Exposure Hierarchy Using Subjective Experience. Participants usually built an exposure hierarchy with clients—a list of OCD triggers ranked based on the intensity of distress they induce—to prepare for the subsequent ERP therapy treatment. Along with initial assessment, participants identified triggers via interviews, self symptom rating tools, and the self symptom monitoring homework (T1-T5, T7) and rated the intensity of each trigger. Most participants adopted the Subjective Units of Distress Scale (SUDS) (Miegel et al., 2022), a 100-point Likert scale (T1), for self distress assessment. Clients were asked to assign each trigger a SUDS score to finalize the exposure hierarchy. Uniquely, T3 asked clients to rate their urge to perform compulsions instead of using SUDS to distract them from focusing on the distress associated with triggers, “ If just tracking SUDS of how anxious you feel now, I mean, that’s what people are doing already. They’re already hyper vigilant about their anxiety. So I need them thinking about something else other than how anxious they feel.”
Furthermore, similar to our findings in Study 1, some participants examined the properties of triggers that can make a certain trigger less or more intense to build a finer-grained hierarchy (T2, T3, T5), as T2 elucidated, “ There is continuing assessment that’s happening during that hierarchy building. Because then I’m asking questions like, what would make [this trigger] easier. What would make it harder? Maybe they’re just telling me like, ‘well… it depends on if someone’s with me, or if I’m by myself; it depends on how many floors there are in the building…’ whatever variables are important to the exposure.”
Designing Diverse Exposures. Based on the exposure hierarchy, participants designed various types of exposures, including in vivo exposures (T1-T5, T7), imaginal exposures (T2-T7), and material exposures (T2, T3, T5, T6). In an in vivo exposure, the client is expected to confront the trigger in real world while supervised by the therapist. For exposures that are difficult to create in real world, imaginal exposures were used where participants created a script together with the client and guide them to imagine a scenario. Although imaginal exposure is less triggering, it can be effective when used appropriately. Two participants (T2, T3) used ChatGPT (OpenAI, 2024) to generate scripts for imaginal exposure. For instance, T2 wrote an obituary using ChatGPT for a client who worried about death. In addition, material exposures, such as videos, images of triggers, and VR exposure, were typically used to bridge the gap between an imaginal exposure and in vivo exposure.
Exposure Practices. Participants guided clients to participate in exposure practices based on the exposure hierarchy, starting from low intensity triggers. Before clients are exposed to triggers, participants first teach them strategies to tolerate anxiety without doing compulsions, aligning with our findings in Study 1. During the exposure task, participants would continue to monitor clients’ SUDS level or urge for compulsions (T3) by periodic polling. Typically participants ended exposure practices when clients’ SUDS decreases (T2, T4, T7).
Besides using SUDS, participants detected clients’ distress level by observing their physical behaviors (T3, T7), such as fidgeting, head shaking, tight chest, staring off the trigger, or sometimes an outward expression of distress. Whenever these happened, participants brought them up in the discussion after the exposure practice to learn what happened. In addition to inhibiting cognitive compulsions, participants also identified clients’ engagement of mental compulsion during the exposure by rating their mental effort (T2).
During high intensity exposure practices where clients could be too afraid to continue, participants would dynamically adjust the exposure intensity (T1-T7). Some participants allowed milder compulsions compared to clients’ usual routines, such as compulsions with shorter duration (e.g., washing hands for shorter time, T3, T7) and lower frequency (e.g., less frequent checking, T2). Additionally, participants decomposed the exposure task into simpler steps (T1, T4). For example, T1 demonstrated an example where a client avoided opening a new book due to fear of contaminating the book with fingerprints. Thus instead of asking them to open the book right away, the client was asked to start with a simpler exposure task—taking the book off the shelf.

### v2

@@ 3. Method
Table 2. Demographic information and professional background of the seven OCD professionals.
3.1. Participants
Our study involved two key stakeholders in the OCD domain—individuals who experience OCD and professionals who specialize in OCD treatment—to develop a comprehensive understanding of the current practices and challenges, ensuring that our design insights would integrate both lived experience of people with OCD and professional recommendations.
3.1.1. People with OCD
We recruited 10 participants with OCD (7 female and 3 male) whose ages ranged from 19 to 47 ($Mean$ = 28.5, $SD$ = 8.5). Participants’ demographic information and OCD types are detailed in Table 1. While our participant gender coverage skewed towards women, it was consistent with the higher prevalence of OCD in female adults than male adults (Mathes et al., 2019). We recruited participants from online OCD support groups and the research email service at our university. A participant was eligible if they were (1) at least 18 years old, (2) diagnosed with mild to moderate OCD, and (3) were willing to disclose their OCD symptoms with us. We did not recruit people experiencing severe OCD at the time of the study because our interview questions can be triggering without adequate ability to manage symptoms independently. Participants were compensated for \$20 per hour.
Our participants demonstrated a wide range of OCD types (Table 1), covering most common obsessions and compulsions in the Y-BOCS Symptom Checklist (Goodman et al., 1989). Six participants (P1, P5, P6-9) had responsibility OCD, whose obsessions were around not being responsible enough or fear of making mistakes; their common compulsions included checking if the door is properly locked. Five participants (P1, P3, P6-P8) had contamination OCD (e.g., dirt, germ) with common compulsions being excessive hand washing and cleaning. Two (P4, P10) had just-right OCD who were obsessed with symmetry or specific position of objects with corresponding compulsions of fixing asymmetry (P4) and objects that are not where they should be (P10). P2 experienced harm OCD, having intrusive thoughts about harm occurring to herself and people around her with compulsion of ruminating about the negative thoughts and avoiding the triggers. She also experienced sex-related obsessive thoughts (sexual OCD) when interacting with men and avoided looking at them to ease her distress. Consistent with the distribution of different OCD subtypes (Ball et al., 1996), more than half of our participants had contamination OCD and responsibility OCD. Table A.1 detailed participants’ OCD subtypes.
3.1.2. OCD Professionals
We recruited seven OCD professionals (5 female, 2 male). Their ages ranged from 29 to 60 ($Mean$ = 43.9, $SD$ = 10.4). All participants had extensive experience working as mental health providers (at least two years), with four (T1, T2, T4, T6) having more than 10 years of experiences. Participants had experiences in treating OCD across various severity levels (from mild to severe) and in different practice modalities (both hybrid and virtual). Table 2 shows participants’ detailed information.
We recruit OCD professionals by sending emails to local OCD therapists listed on the International OCD Foundation, a nonprofit organization providing resources and education for people with OCD (IOCDF, 2024). Participants were eligible if they were over 18 and had experience conducting OCD related therapy. We limited the recruitment to individuals who spoke English and located in the United States. Each participant received \$40/hour as compensation.
3.2. Procedure
We conducted single-session semi-structured interviews with both participants with OCD and OCD professionals via Zoom. Each session lasted for about 90 minutes. Our interview focused on different perspectives between the OCD individuals and the professionals.
For individuals with OCD, we focused on their unique triggers, experiences, challenges, and coping strategies in daily life. Specifically, we first asked about their demographic information (e.g., age, gender identity) and diagnosis of OCD. Then we asked participants about their detailed personal experience with OCD—we prompted them about their obsessions, triggers, and compulsions in different living environments (e.g., home, workplace, public spaces). For triggers, we specifically asked for fine-grained details with a goal of identifying the characteristics and rationales of the triggers. We then asked participants about their strategies and technology use to manage their OCD symptoms in daily life, either recommended by therapists or developed by themselves. We also asked about their therapy experiences and use of other resources to assist with symptom management. Finally, participants discussed ideas about desired real-life interventions for symptom self-management.
For OCD professionals, we focused on their strategies and practices in OCD therapy sessions to potentially inspire intervention design. We first asked about professionals’ demographic information, certifications, and experience as a mental health therapy provider. We then asked about the types of therapy they conducted, the overall therapy process (e.g., treatment stages), the strategies and technologies used at different treatment stages, and how they evaluated client progress. In addition to in-session practices, we asked therapists about out-of-session strategies they recommend, such as assigned homework and external support. We then had them discuss their current challenges in therapy practices. Lastly, we asked therapists about their familiarity with and opinions on emerging technologies, such as phone-based self-administered therapy. When therapists were unfamiliar with certain technologies (e.g., virtual reality), we provided a brief, neutral description to ensure informed responses and minimize misconceptions.
3.3. Analysis
We audio recorded and transcribed all interviews via an online transcription service (otter.ai [^1] https://otter.ai). One researcher manually cleaned all transcripts and corrected transcribing errors by checking the recorded videos. We analyzed the data using thematic analysis (Clarke et al., 2015; Braun and Clarke, 2006). We selected two representative transcripts from OCD participants and two from OCD professionals as samples to develop initial codebooks. Two researchers coded the samples independently with open coding. Then, they discussed and reconciled their codes to resolve any differences and developed two initial codebooks upon agreement, one for participants with OCD and one for OCD professionals. One researcher then coded the remaining transcripts from the two groups of participants based on the codebooks, respectively. If a new code was constructed, the two researchers discussed on the code and updated the codebooks upon agreement.
We first derived themes for OCD participants and OCD professionals separately, where we used a hybrid of inductive and deductive approaches, a method commonly adopted in qualitative research (Oewel et al., 2024; Swain, 2018; Braun and Clarke, 2006). We started with high-level themes based on fundamental concepts of OCD triggers, compulsions, therapy treatment, and self-management identified in prior work (Wu and Lewin, 2017; Knapton, 2016; Blakey et al., 2017; Wu and Lewin, 2017; Yule and Whittal, 2017) using a deductive approach (Crabtree and Miller, 1992). We then used an inductive approach (Boyatzis, 1998) to generate sub-themes and identify new themes by grouping related codes via affinity diagramming (Clarke et al., 2015; Braun and Clarke, 2006) for each participant group. Finally, we generated final themes and sub-themes across participants by comparing and combining themes from the two groups. Two researchers cross-referenced the original data, the two codebooks, and themes to make final adjustment, ensuring all codes were grouped in the correct themes. Our analysis resulted in over 700 codes and 20 themes and sub-themes.
3.4. Ethical Considerations & Reflexivity Statement
Our interviews involved discussing participants’ OCD symptoms, which could cause mental distress. To provide a safe and comfortable environment for OCD participants, we consulted a clinical psychologist (a co-author) with experience treating OCD to ensure appropriate language was used in the interview. OCD participants were informed that they could skip any uncomfortable questions and turn off their camera if preferred, and they could exit the study anytime when needed.
We recognize that researchers’ identities and background may influence our data interpretation. The primary researcher identifies as a person with OCD. His background and lived experience with OCD enabled us to develop appropriate questioning routes, better empathize with participants, and interpret their responses via a unique lens. Moreover, the research team comprises diverse expertise, including HCI, accessibility, intelligent system design, and OCD therapy, enabling us to better contextualize the study data in existing technology and mental health literature, thus deriving suitable technology implications to support people with OCD.
We interviewed both people with OCD and therapists who treat OCD to gather diverse perspectives on symptom self-management strategies and needs. While therapists offer insights grounded in treatment principles and their experience with clients, we acknowledge their positionality as mental health professionals rather than individuals with lived experience. Their interpretations may not fully capture, and may at times diverge from, how people with OCD understand and relate to their experiences. As such, our analysis prioritized the narratives of people with OCD and used therapist perspectives as a complementary lens. When tensions between the two groups emerged, we present them transparently without privileging therapists’ views.
4. Findings
Our study findings go beyond the conventional clinical view of OCD and takes a unique sociotechnial lens to understand the OCD individuals’ unique triggers, compulsions, and coping strategies as well as highlighting the overlaps and contrast between the perspectives of people with OCD and the professionals. We detail our findings below.
4.1. Redefining OCD Triggers
We first investigated the OCD triggers to determine what triggers to recognize and how to recognize triggers for just-in-time interventions for OCD. We highlight that triggers go beyond “a common list” followed by prior work (Mataix-Cols et al., 2009) but include various types in different forms. Moreover, it is not the trigger type (e.g., a trashcan) in itself but rather its properties (e.g., fullness of the trashcan) that can significantly affect an OCD individual’s experiences.
4.1.1. Triggers can be anything
Unlike prior work that limits OCD triggers to a set of ”common triggers” based on basic OCD themes (e.g., contamination, harm) (Mataix-Cols et al., 2009), our findings show that OCD triggers are highly individualized. While most triggers reported by OCD participants are specific physical objects (i.e., triggering objects), such as trash cans for contamination OCD (P1) or knives for harm OCD (P2), participants also revealed less recognized trigger forms, including contexts, sounds, and even thoughts.
We found that OCD can be triggered by participants’ physical context, such as specific locations or times (P1, P3, P7-9). For instance, participants with contamination OCD (P1, P3, P8) were triggered every time they returned home, since germs brought from the outside can contaminate their home environment. Interestingly, weather can also be a trigger. For example, P7 indicated her fear of hitting a pedestrian emerged with low visibility on the road during rain or snow. OCD triggers can also be audible (P3, P6-8, P10). For example, coughing sound can immediately trigger P3’s contamination OCD, and accident-related keywords in a conversation would trigger P8 to check his own apartment for safety. In addition, some participants reported mental triggers (P5), such as unspecified worries, without clear links to physical objects. Moreover, OCD episodes could sometimes occur even without identifiable triggers (P1, P4, P6, P10).
Inspiration:
Inspiration: The diverse forms of triggers reveal the difficulty of real-time trigger recognition and highlight the needs for the use of multi-modal sensory data in trigger identification.
4.1.2. Trigger Intensity: Properties over Objects
Beyond the trigger type, the intensity of triggers (i.e., the mental distress induced by a certain trigger) plays an important role in determining whether an individual would be triggered. Our OCD professionals emphasized the importance of asking clients not just about the triggers themselves, but also about the factors that influence their intensity (T2, T3, T5). Specifically, the OCD professionals often work with the clients to create a trigger hierarchy—a ranked list of OCD triggers ordered based on their intensity. However, this process usually follows a specific template by asking about clients’ experience with a checklist of “common triggers” (e.g., Y-BOCS (Goodman et al., 1989)) and then asking them to rate their distress caused by a specific trigger (T1-T2, T4-T7) or their urge to performance compulsion (T3). While some professionals allowed their clients to add more diverse triggers by logging their encounters outside the therapy session (T1–T5, T7), we found that there was a lack of comprehensive method in clinic to identify all possible triggers. Moreover, the method of evaluating trigger intensity was completely based on client’s subject ratings, without systematic characterization of the factors that affect trigger intensity.
Despite the lack of clinical standard, our study data from OCD individuals revealed that, for physical triggers, instead of the object type in itself, it was often the specific properties of the object that determined the trigger intensity, leading to varying levels of anxiety and compulsions. Below, we summarize key properties that affect trigger intensity from the experiences of different participants with OCD.
Sensory Attributes. Participants reported multiple sensory attributes that can alter the intensity of a trigger. Visual attributes such as color can affect people’s perception of a trigger (P3, P7). For example, P7’s contamination OCD were triggered when the dirt presents high contrast against its background, such as coffee grind on a white countertop. While the same amount of white salt (also considered as dirt) with low contrast to the countertop would not be as triggering. In addition to color, some specific shapes can also be more triggering than others, such as sharp edges. P2 felt more uncomfortable when interacting with objects that have sharp tips in her home environment, including knives with sharp blade, pencils with sharp tips, and the newel posts with sharp corners on a staircase. Moreover, tactile attributes of an object could be another factor that affected trigger intensity (P1, P7). For instance, the sticky texture of sauce (e.g., ketchup) caused more contamination concerns to P1 and P7 than solid food.
Affordances. We found triggering objects that afforded direct contact with the user can cause higher level of distress to participants with contamination OCD (P1, P3, P6-8). For example, door handles in public spaces and handrails in transportation are common triggers because it is meant to be touched using hands by many people, thus considered dirtier than surfaces not requiring direct contact (e.g., automatic doors). P1 mentioned trash can as one of her OCD triggers, but the ones with a lid and handle were more triggering for the same aforementioned reason.
Spatial Attributes. Spatial attributes of a triggering object is another property that determines trigger intensity. First, the quantity of triggering objects can affect the intensity of contamination OCD. P3 tended to avoid occasions with big crowds of people more than others due to concerns about contagious diseases. Second, the proximity of a trigger to a person with OCD is another factor that alters trigger intensity. According to P2, a knife in hand caused more harm-related intrusive thoughts to her than a knife on the counter top. Interestingly, trigger intensity can depend on the pose and orientation of a physical trigger. P2’s sexual intrusive thoughts were more likely triggered when a man was in a position that “resembled those in her intrusive thoughts.”
Familiarity. Trigger intensity varied depending on participants’ familiarity with their triggers. For participants with harm OCD (P2), familiar triggers can be more intense. For example, P2 experienced more intrusive thoughts about accidentally hurting people when around family or friends rather than strangers. Conversely, participants with contamination OCD (P1, P8) found familiar triggers less intense. For instance, P1 was not distressed by food residue if she knew who was responsible for it. “Right now, in my kitchen, there’s a few dirty dishes in the sink, and there is a pot on the stove, and a box of pasta that needs to be put away, but I know that my roommate made Pasta last night, and so I’m not at all stressed. Because I know exactly where everything came from.” Similarly, P8 was more concerned about germs brought into his apartment by strangers than by his own family, whose hygiene he trusted.
Internal Factors Affecting Trigger Intensity. In addition to trigger properties, participants with OCD highlighted several internal factors that influenced trigger intensity. Six participants (P1-4, P6, P7) further pointed out that their OCD triggers intensified when experiencing mental stress. It can be from schoolwork (P1, P2, P4), being in unfamiliar environments (P1) and personal life issues (P4). Additionally, poor self-care, such as sleep deprivation and physical inactivity, also exacerbated trigger intensity (P1, P2). Interestingly, unlike mental stress, physical stress could sometimes help reduce trigger intensity (P7, P8). As P8 explained, “if I’m pretty tired, my brain can’t put so much energy towards the OCD.” Similarly, trigger intensity can be reduced when short-term memory is fully occupied by other tasks (P2, P6), e.g., when working (P6).
Inspiration:
Inspiration: Trigger intensity is affected by not only trigger itself, but also properties of triggers whose variation can result in significantly different anxiety levels. As a result, trigger recognition for just-in-time intervention should not only consider object-level detection but also the recognition of certain properties to more accurately predict users’ distress, thus providing suitable intervention.
4.2. How Triggers Shape Compulsions
Besides triggers, compulsion is another important indicator to invoke just-in-time interventions. Timely detection of compulsive behaviors in daily life is crucial for delivering effective interventions. We thus investigated the diverse compulsive behaviors and their relationships to triggers to inspire compulsion detection.
4.2.1. Distinguishing Compulsive Behaviors
Compulsive behavior is harder to identify than triggers since compulsions often manifest in ways people with OCD might not even realize, such as avoidance (e.g., avoiding contact with objects deemed dirty by people with contamination OCD) and mental compulsion (e.g., rumination) (T1-T3, T5-T7). Therefore, in addition to identifying the typical compulsions (e.g., excessive hand washing) using standardized checklist (e.g., Y-BOCS (Goodman et al., 1989)), professionals would also ask clients about the presence of covert compulsions directly, such as avoidance and mental compulsions (T1-T3, T5-T7). In general, our professionals outlined three criteria to help distinguish compulsive behaviors from regular behaviors and assess their impact on daily life:
@@ 4.2.1. Distinguishing Compulsive Behaviors
Inspiration:
Inspiration: These compulsion identification strategies from professionals could potentially be adapted to support compulsive behavior recognition in real life, for example, with the help of machine learning models that can distinguish common sense from irregular behaviors (Aggarwal et al., 2021).
4.2.2. Connecting Compulsions to Triggers
Similar to OCD triggers, our OCD participants experienced diverse compulsive behaviors. Unlike prior OCD literature that focused on recognizing compulsive behaviors in itself (e.g., washing hands) (Starcevic et al., 2011; Goodman et al., 1989), we found that our participants constantly associated their compulsive behaviors with triggers. Understanding the rationale behind compulsions, particularly how individuals respond to specific triggers, is crucial in interpreting compulsive behaviors, which can potentially help enhance compulsion detection and facilitate intervention delivery at the right timing. We thus categorize compulsions into five categories based on their relations to triggers:
Direct Trigger Interaction. Direct trigger interactions are physical behaviors that engage with the trigger to temporarily resolve the distress caused by the trigger. Repeatedly examining the trigger is a common compulsion for participants with responsibility OCD (e.g., checking or jiggling the door locks, P8, P9), and contamination OCD (e.g., checking the color of meat being cooked to ensure doneness, P3). Ordering and organizing is another direct interaction with triggers to precisely control the position, orientation, or order of the triggers to avoid accidents (e.g., keep appliances unplugged when not using, P7, P8), reduce harmful thoughts (e.g., pointing the knife to herself to protect others, P2), and keep life organized and “just right” (e.g., having to park his car in the same spot, P10). Direct trigger interactions can also happen when the OCD trigger is a person’s body part. For example, P4 would pull her eyelash or hair so both sides look even whenever she examines her look in the mirror. However, while temporally relieving mental distress, such compulsion could bring long-term mental health issues, “[Hair and eyelash] takes the longest to grow back and it really impacts my confidence and stuff, because it’s noticeable.” (P4)
Trigger Avoidance. Trigger avoidance is another type of compulsions that minimize the interaction with triggers (P1, P3, P6, P8), We found that the extent of avoidance was a spectrum, ranging from partial avoidance, such as rolling their hands into sleeves (P3, P8), or using as few fingers as possible (P6) when touching door knobs, to complete avoidance, such as avoiding eye contact with the trigger (P2). Trigger avoidance can lead to severe functional impairment, threatening OCD people’s quality of life. P8 expressed his struggle to engage in essential daily activities due to his contamination avoidance, “If there are certain activities that I think are dangerous that’s not safe or something bad could happen. It’s gonna get in my head… The biggest challenges are just refusing to engage in a lot of activities to avoid… And I’m not gonna be able to fix the issue the way I want. And I’m gonna be stuck somewhere where I can’t do anything about it. ”
Consequence Mitigation. Consequence mitigation is a compulsive behavior that addresses the speculated aftermath after interacting with an object trigger when it is inevitable. For participants with contamination OCD, changing/cleaning clothes when coming back home and washing body parts after contacting something ‘dirty’ are common compulsions. Additionally, P3 mentioned self-body scanning for potential symptoms when encountering someone appearing sick. For participants with responsibility OCD, repeated checking can also be consequence-oriented. For example, participants who worried about hitting pedestrians while driving would check for marks on the car later (P7) or look at rear mirror to confirm no person was run over (P9). P7 described this lengthy process, “It used to take me like 10 minutes to check all of my car because I would get out and I would check like the doors and the windows, and I would check the hood of the car and the back of the car for scratches.”
Mental Compulsions. In addition to physical compulsive behaviors, seven participants reported experiencing mental compulsions (P2-5, P7, P9, P10) where they attempted to neutralize the influence of triggers internally. For participants with just-right OCD, a common mental compulsion was internal counting, for instance, internally counting delayed time when their schedule was disrupted (P10). Other mental compulsions included self-questioning rumination, such as convincing oneself of being a good person despite having intrusive thoughts (P2, P5), catastrophizing about consequences of disrupted routines (P10), and self-reassurance to suppress negative thoughts (P7).
Inspiration:
Inspiration: Connecting triggers and context information to compulsions can potentially facilitate compulsion recognition by tracking not only user behaviors but also how specific triggers shape the compulsive reaction of the user. As such, a user’s compulsion patterns can be accurately characterized, allowing for more contextualized and personalized intervention design.
4.3. Managing OCD at Different Stages
To manage their OCD symptoms, our OCD participants adopted a range of strategies learned from OCD therapy (P1–5, P7, P9, P10), support groups (P5), online educational resources (P8), and personal experiences (P6, P10). These strategies were applied at different stages of an OCD episode—from the onset of obsessive thoughts to the execution of compulsions. By integrating perspectives from both people with OCD and professionals, we observed substantial overlap in strategy use, but also noted some inconsistencies in how these strategies were applied or recommended. We summarize these coping strategies to inform the feedback design of just-in-time interventions for OCD self-management technologies.
4.3.1. Accepting & Challenging Obsessive Thoughts
OCD participants employed different mental strategies to address their obsessions, adapted from ACT and cognitive therapy (see Section 2.1).
Three participants with OCD and nearly all professionals (except T4) used or recommended ACT-based strategies to help shift focus from obsessions to value-driven actions. These strategies included: (1) Thought Diffusion (P3, P5, T1, T5)—recognizing and naming the intrusive thought to create psychological distance from the thought; (2) Embracing Uncertainty (P2, P3, T5-T7)—using self-talk like “maybe, maybe not” to accept uncertain outcomes; (3) Refocusing on Values (P5, T1-T3, T5)—identifying values (i.e., important aspects of life) and committing to behaviors aligned with those values.
While some OCD participants benefited from accepting obsessive thoughts, others managed obsessions by actively challenging and re-evaluating them to reinforce positive thinking (P1, P7), a technique also endorsed by therapists (T3, T5). For example, P7 addressed her intrusive fears of unconsciously hitting a pedestrian while driving by evaluating the evidence: “‘okay, what’s the likelihood that this (hitting pedestrians) actually happened? And you know, wouldn’t someone else have noticed if this had happened? Wouldn’t I have heard, or seen, or touched or smelt things if this had happened?’ So it address like the thought pattern.”
Inspiration:
Inspiration: Leveraging the self-reflective nature of these mental strategies, we can design conversation-based persuasive intervention (Orji and Moffatt, 2018) to guide people with OCD to adopt beneficial behaviors and thoughts when obsessions occur, such as intelligent agents that help users navigate their thought processes and address cognitive pitfalls.
4.3.2. Compulsion Prevention via Real Life Exposure Practices
When obsessive thoughts were too intense to address or dismiss, participants with OCD resorted to strategies learned from ERP (P1-3, P5, P7, P10) to effectively train their ability to tolerate the anxiety without performing compulsions.
Planned vs. Natural Exposure Practice. To improve symptom management ability in a controlled environment, participants would do planned exposure practice as instructed by their therapists (P1-3, P5, P7, P10). In a typical planned exposure practice, people with OCD expose themselves to a self-created triggering environment starting from triggers with moderate intensity on the trigger hierarchy at a scheduled time, and resist compulsions. P3 articulated her exposure practice targeting the fear of touching batteries, “It (exposure practice) was holding a battery in my hand, rubbing the battery on my arms, putting it in my pocket, rubbing it on my clothes, my phone, on my bag… So I, essentially contaminating all these different things [with batteries] and then, not being able to wash my hands or do anything to decontaminate the battery acid… and having to sit with worrying that everything was contaminated.” While tolerating the mental distress, people with OCD were instructed to monitor their anxiety level change (e.g., SUDS level) (T2-T6) during the exposure practice. In addition to self rating, when possible, professionals detected clients’ anxiety level also by observing their physical behaviors (T3, T7), such as fidgeting, head shaking, tight chest, staring off the trigger, or sometimes an outward expression of distress. Typically an exposure practice concludes when their anxiety level decreases (T2, T4, T7).
Although the flexibility of adjusting trigger intensity allowed participants to gradually build up their tolerance to triggers, most triggers were unplanned in real life, and some triggers that induce intrusive thoughts could not be anticipated (P2, T1, T7). Therefore, when confronted with unplanned triggers, participants would seize the opportunity to do a natural exposure practice. Due to the spontaneity of unplanned triggers and unpredictable trigger intensity, natural exposure practices were perceived to be more difficult (P1-P3, P7, P10).
Despite the effectiveness of both planned and natural exposure practices, OCD participants found them challenging especially at the beginning (P2) and tended to avoid exposures when alone (P1). They had to force themselves to complete the practices, since compulsions were too difficult to resist (P1-P3, P10). As P10 described, during an exposure practice where the trigger was overwhelming, “it was mentally painful to try and stop (the compulsion).”
Inspiration:
Inspiration: These findings highlighted the importance of providing timely and context-aware guidance to help users engage with the exposure rather than avoid it, especially at the beginning, when the user is still building confidence in managing their symptoms.
Strategies for Resisting Compulsions. To better support compulsion prevention, OCD participants reported different mental and physical strategies to suppress the urge of doing compulsions. One common strategy is to re-evaluate the necessity of compulsions mentally (P2-4, P6). By reflecting on past OCD episodes, participants reminded themselves that nothing bad happened when compulsions were not done (P2) and that compulsions were never working (P4, T1, T7). Some weighed the harm compulsions caused against the relief they brought (P3, P4). For instance, frequent hand washing killed germs but caused dry, irritated skin (P3). In addition to mental persuasion, participants also employed physical strategies to make compulsions more difficult to act on, e.g., P4 creatively put Vaseline on her eyelash and wore glasses to prevent herself touching her eyelash.
Completely resisting compulsions overnight can be overwhelming. Instead, gradually tapering them—by allowing milder forms of compulsive behavior—can support long-term compulsion reduction. Both OCD participants and professionals in our study employed such strategies. Participants with responsibility OCD used reassurance logging to gradually reduce the intensity and frequency of their compulsions over time (P8, P9). For example, they recorded videos or took photos of their checking behaviors to ease anxiety and avoid the urge to re-check later. As P8 explained: “If I know I’m gonna be gone for a while. I will take a video where I’ll video my entire apartment, and then I’ll also turn the camera around and say, ‘I have checked the faucet this many times; all plugs are unplugged…’ so it’s almost like telling myself through a video in the moment that I’m good to go. And then if I ever get anxious, I can look back at the video.” Similarly, professionals allowed clients to engage in reduced compulsions when full exposure practice felt too distressing. These included compulsive behaviors with shorter duration (e.g., washing hands for a shorter time, T3, T7) and lower frequency (e.g., checking less often, T2).
Inspiration:
Inspiration: In summary, participants effectively reduced the urge for compulsions by drawing on past behaviors and memories of previous OCD episodes. Inspired by these strategies, symptom and behavior logging could be leveraged to design just-in-time intervention methods that increase the cognitive and physical burden of engaging in compulsions. Tapering strategy can also be translated into just-in-time interventions by prompting users with milder alternatives to their compulsions, helping them gradually build tolerance to anxiety.